1. What is the molarity of a solution that contains 0.5 moles of HCl in 200ml of water? 
  2. What is the molarity of a solution that has 5g of H2SO4 in 150ml of water?
  3. 15g of Sodium Chloride in 1.5L of water?
  4. 22g of magnesium chloride in 750ml of water?
  5. 2.7x 10-4g of sodium chloride in 15mL of water?



pH problems


  1. What is the pH of a 0.2M HCl solution?
  2. Find the pH of a 0.0015M solution of nitric acid.
  3. 0.0025M sodium hydroxide solution
  4. Find the pH when 5g of HCl is dissolved in 150ml of water
  5. Find the pH when 4g of NaOH is added to 500ml of water
1
Expert's answer
2021-06-02T06:31:23-0400

1)"Molarity= moles \/litre"litre

Molarity = 0.5/1000

=0.5× 10-3M

2)5g of H2SO4= 5/98 = 0.051moles

Molarity= 0.051/1000= 5.1× 10-5M

3)15g of Sodium Chloride

= 15/58.4= 0.256 moles

Molarity=( 0.256× 1.5)1000

= 3.84×10-3 M

4)22g of magnesium chloride

=22/95 = 0.232moles

Molarity=(0.232×0.75)/1000

=1.74×10-4M

5)2.7x 10-4g of sodium chloride

=2.7× 10-4/58.4 =4.6× 10-6moles

Molarity=( 4.6× 10-6×0.15)/1000

= 6.9×10-10M